OpenStudy (anonymous):

For f(x) = x^1/3, f(x) is not differentiable at x=___? hint: try to find the derivative by using the definition

myininaya (myininaya):

\[f'(x)=\frac{1}{3}x^{\frac{1}{3}-1}=\frac{1}{3x^\frac{2}{3}}\] when is the bottom zero?

OpenStudy (anonymous):

when x=0?